# Landau-Zener-Stückelberg interference in edge state pumping

### v1

@@ Abstract
The adiabatic edge state pumping (ESP) in one dimensional model, which has important applications in topological phase transition and the associated implementation of edge states in quantum simulation, has been widely performed in both theories and experiments. This phenomenon has been verified in some small physical models, yet some fundamental issues about this process have not been clarified. In this paper, we revisit this problem of ESP and pinpoint a pair of non-adiabatic points in the band levels, at which the adiabatic condition breaks down. We determine the two points using the criteria of non-adiabaticity. As a result, the oscillation of ESP as evolution time varies can be resolved in terms of Landau-Zener-Stückelberg (LZS) interference. Furthermore, in the presence of disorder, we show that the ESP may break down for the anticrossing between the edge and the bulk levels, where the non-adiabaticity diverges. Thus in a relatively long chain with weak disorder, we demonstrate the failure of the ESP. This new type of ESP unveiled in this work is readily accessible in experiment, and shall therefore lead to a down-to-earth platform for the intriguing LZS dynamics.
@@
Introduction: The discovery of topological matters have stimulated widespread investigations of their edge states ensured by bulk-edge correspondence in quantum simulation Qi and Zhang (2011); Hasan and Kane (2010); Schnyder et al. (2008); Ozawa et al. (2019); Georgescu et al. (2014). In one dimensional (1D) topological models, the edge states are localized at the open ends, which can gradually transfer from one end to the other by adiabatically tuning the parameters, termed as Thouless (or adiabatic) pumping Thouless (1983); Citro and Aidelsburger (2023); Aleiner and Andreev (1998); Sun et al. (2022); Jürgensen et al. (2021, 2023); Cerjan et al. (2020). It has been believed that when the pumping process is slow enough, this kind of adiabatic process can always be achieved. Following this tenet, the adiabatic pumping in the 1D chains with quasiperiodic potentials were experimentally investigated using coupled single-mode waveguides Lahini et al. (2009); Kraus et al. (2012); Rechtsman et al. (2013); Wang et al. (2019); Jürgensen et al. (2021). This idea was also generalized to higher dimensions in terms of synthetic dimensions Yuan et al. (2018); Zilberberg et al. (2018); Yang et al. (2020). Since the edge states are protected by topology, this adiabatic pumping withstands weak disorders Nakajima et al. (2021); Kolodrubetz et al. (2018) and nonlinear interactions Jürgensen et al. (2021); Tangpanitanon et al. (2016); Mostaan et al. (2022); Jürgensen et al. (2023); Qian et al. (2011). In recent years, this adiabatic pumping has been widely examined using electromechanical patches Xia et al. (2021); Wang et al. (2023a); Rosa et al. (2019), acoustic waveguides Apigo et al. (2019); Rosa et al. (2019); Chen et al. (2021); You et al. (2022); Hu et al. (2024), photonic crystals Ke et al. (2016); Meier et al. (2016); Cheng et al. (2022); Adiyatullin et al. (2023); Poshakinskiy et al. (2014), and even in the Floquet-Bloch band Upreti et al. (2020); Minguzzi et al. (2022); Adiyatullin et al. (2023). The chain sizes for all systems above range about 9 - 100. The related adiabatic transport of matter in terms of Thouless pumping even with many-body interaction was also explored in experiments Schweizer et al. (2016); Viebahn et al. (2024); You et al. (2022) for applications in quantum matter transport, state transfer and information processing.
@@
This edge state pumping (ESP) is fundamentally different from the adiabatic transition in the two-level system Cohen et al. (2019); Berry (1984), in which the adiabaticity is guaranteed by the finite gap between the two levels. In the continuum models, the number of energy levels is determined by the lattice size $L$, in which the large $L$ is the smaller the level spacing $\Delta$ will be. Thus there are two competitive mechanisms to influence the adiabaticity during the pumping process: the overlap between the edge state and its nearest neighboring bulk state, and the energy gap between them, both of which decrease with the increasing of $L$. Therefore when the former term decreases faster than the latter, the adiabatic condition can be fulfilled; otherwise, it becomes possible that the adiabatic condition may break down, yielding non-adiabatic evolution. Therefore, while the ESP has been verified in experiment Kraus et al. (2012); Rechtsman et al. (2013); Switkes et al. (1999); Lahini et al. (2009); Rosa et al. (2019); Verbin et al. (2015, 2013); Zilberberg et al. (2018); Yang et al. (2020); Nakajima et al. (2021); Kolodrubetz et al. (2018); Jürgensen et al. (2021); Tangpanitanon et al. (2016); Mostaan et al. (2022); You et al. (2022); Jürgensen et al. (2023); Xia et al. (2021); Wang et al. (2023a); Apigo et al. (2019); Chen et al. (2021); Hu et al. (2024); Ke et al. (2016); Meier et al. (2016); Cheng et al. (2022); Adiyatullin et al. (2023); Xiang et al. (2023); Rajagopal et al. (2019); Cerjan et al. (2020); Riva et al. (2020), the stringent validity of adiabatic condition, which is essential for ESP, remains yet to be verified. Furthermore, the disorder effect in ESP also needs to be clarified Rajagopal et al. (2019); Yu et al. (2020); Cerjan et al. (2020); Tangpanitanon et al. (2016). These two issues will affect ESP severely in the large $L$ limit with vanished coupling between the edge states Chen et al. (2021).
@@
We then revisit this problem and present a new picture for the ESP in these models. (I) We show that there are two non-adiabatic points (NAPs) during the ESP process, at which the adiabaticity breaks down even in the slowly-varying limit. This is when the ESP process reduces to the Landau-Zener-Stückelberg (LZS) interference Berry (2009); Shevchenko et al. (2010); Wang et al. (2015); Boyers et al. (2020); Chen et al. (2021); Qiao et al. (2023), resulting in transient oscillation of ESP as evolution time $T$ changes. (II) In the presence of weak disorder when the localization length $\xi$ is larger than the system size $L$, the anticrossing of the bulk states and edge state breaks adiabaticity when the edge state is immersed into the (continuum) bulk states Lahini et al. (2009). In this specific case, the ESP will cease to happen. These new mechanisms can be meticulously examined in simulating platforms with ultracold atoms Schomerus (2013); Viebahn et al. (2019); Nakajima et al. (2021), optical and acoustic waveguides Xia et al. (2021); Wang et al. (2023a) and electric circuits Li et al. (2023); Mei et al. (2018), etc..
@@
with incommensurate potential $V_{n}=V[1+\lambda\cos(\pi\alpha n+\theta)]$, where $V$ is the hopping strength between neighboring sites, $\alpha$ is the irrational constant and $\hat{c}_{n}$ ($\hat{c}_{n}^{\dagger}$) is the annihilation (creation) operator at site $n$ [see Fig. 1 (a)]. Hereafter, we use $V=8/15$, $\lambda=0.6$, $\alpha=(\sqrt{5}+1)/2$ (parameters from Ref. Kraus et al. (2012)), with uniform disorder $\xi_{n}\in[-1/2,1/2]$. The parameter $\theta$ is used as the adiabatic parameter, by which the edge state transfers from the left end to the right within a finite pump time $T$. The energy levels in this model as a function of the phase $\theta$ are presented in Fig. 1 (b), in which the two thick solid blue lines represent the edge modes, respectively, with opposite energies. We focus on the pumping of edge state with $E>0$ in this paper. To this end, we expand the wave function using $|\psi(t)\rangle=\sum_{n}c_{n}(t)|n\rangle$, where $|n\rangle=\hat{c}_{n}^{\dagger}|0\rangle$, and the coefficient $c_{n}$ is determined by Schrodinger equation $i\partial_{t}|\psi(t)\rangle=\hat{H}(t)|\psi(t)\rangle$, with $\theta=\theta(t)$ to be the time-dependent tuning parameter.
@@
Figure 2: Distribution of wave function density $|\psi(x_{i},t)|^{2}$ in spacetime and its mean trajectories $\bar{X}(t)=\sum_{i}x_{i}|\psi(x_{i},t)|^{2}$ (green lines) for three pump times: (a) $T=1300$, (b) $T=1750$ and (c) $T=9500$, with $L=42$ and without disorder. (d) Occupation probability $\rho_{n}$ versus pump time $T$ for edge state $|\psi_{26}\rangle$ and bulk state $|\psi_{27}\rangle$. Note that the three pump times are chosen from three extrema on $\rho_{26}$. Inset shows the optimum pump time $T^{*}$ for the first maximum [in blue asterisk in panel (d)] versus the chain length $L$.
@@
Our central idea is that during the ESP, there are two points which are most vulnerable to break the adiabatic condition. This can be understood as follows. When the edge state is fully localized at the end, the overlap between the edge states and the bulk state induced by variation of Hamiltonian $\delta H$, which may be defined as $O\sim{|\langle\psi_{\rm e}(\theta)|\delta H|\psi_{\rm b}(\theta)\rangle|}$ with $|\psi_{\rm e}\rangle$ for the edge state and $|\psi_{\rm b}\rangle$ the bulk state respectively, will decay in a way of $1/L$. Meanwhile, the energy difference $\delta E$ between the bulk and edge states will approach a constant; see regime A in Fig. 1 (b). Furthermore, for $\theta$ in regime B in Fig. 1 (b) when the edge states become extended, the overlap between them $O\sim 1/L$, while the energy gap between these two states also scales as $\delta E\sim 1/L$. Under these two conditions the the adiabatic condition can be easily fulfilled by controlling the sweeping rate even in the large $L$ limit Grabarits et al. (2024). The adiabaticity may break down between the regimes A and B with the increasing of system size $L$, which are marked by arrows as NAPs.
@@
Thus when the wave functions is restricted to occur between the edge state and the nearest neighboring bulk state and by projecting the wave function to these two states, the dynamics will be reduced to the LZS interference process. To gain some insight into this process, let us first consider the following two-level model [Fig. 1 (c)]
@@
where $\hat{\sigma}_{x}$ and $\hat{\sigma}_{z}$ are Pauli matrices, $g$, $A$ are tunneling amplitude and bias amplitude respectively, and $T$ is the total evolution time. The Landau-Zener tunneling happens at $t=T/4$ and $T=3T/4$. In Fig. 1 (d) the dynamics of wave functions is presented, showing that by slightly changing $T$, the final state may be significantly different. The interference depends strongly on the relative phase $\phi$ between the two levels [Fig. 1 (c)], which has been widely verified in experiments Huang et al. (2011); Quintana et al. (2013); Korkusinski et al. (2017); Wang et al. (2023b).
ESP without disorder: We first consider the case without disorder ($W=0$). Initially, we assume the wave function to be the edge state $|\psi_{n}(\theta)\rangle$ and define
@@
where $|\psi_{n}(\theta)\rangle$ is the $n$-th instantaneous eigenstate with phase $\theta(t)$ and $|\psi(t)\rangle$ is the time-dependent wave function. The results for different $T$ are presented in Fig. 2 (a) - (c) for size $L=42$, and $n=26$ as the index of the edge state. For $T=1300$ and $9500$, the left end edge state can be transferred to the right end. However, when $T=1750$, the state will not evolve to the right end. This result is in stark contrast to adiabatic theory in which the larger $T$ is the better the adiabaticity will be Berry (1984). In Fig. 2 (d), the oscillation of $\rho_{26}$ and $\rho_{27}$ versus pumping time $T$, is a signature of strong nonadiabaticity before entering the adiabatic limit. Furthermore, $\rho_{26}+\rho_{27}\simeq 1$, indicating that only these two states are important during evolution, justifying that the dynamics can be reduced to the two-level model for LZS process. This oscillation was found numerically in Ref. Longhi (2019), but with its fundamental origin not revealed. The back and forth of the mean wave function position can be readily measured in the current experiment. In the inset of (d), we plot the the time $T^{*}$ for the first maximum ESP (in blue asterisk) as a function of system length $L$, showing that $T^{*}\propto L$, indicating that the longer the chain is, the longer the required time $T^{*}$ for ESP will be.
To clarify the LZS dynamics, we plot the evolution of $\rho_{n}$ ($n=26-29$) for $L=42$ as a function of time for three typical $T$ in Fig. 3 (a) - (c). In (a) at the first NAP, $\rho_{26}$ drops suddenly to about 0.5, indicating that the two states are almost equally occupied. However at the second NAP, the probability will be transferred back to unit, indicating that the bulk states of $|\psi_{27}\rangle$ will be transferred to the edge state $|\psi_{26}\rangle$. This is a surprising example to realise a perfect ESP yet with non-adiabatic process. In (b) with even more pumping time $T=1750$, we find that after the two NAPs, the edge state will finally be transferred to the bulk state, with $\rho_{26}\sim 0$. In (c), the adiabatic ESP can be achieved when $T$ is large enough. These behaviors are qualitatively the same as the LZS process in Fig. 1 (d) by varying of $T$.
Figure 3: Evolution of $\rho(t)$ for the edge state $|\psi_{26}\rangle$ and the three nearest bulk states $|\psi_{27}\rangle$, $|\psi_{28}\rangle$ and $|\psi_{29}\rangle$ for pump time (a) $T=1300$, (b) $T=1750$ and (c) $T=9500$ (Here, $L=42$). (d) non-adiabaticity ${N}_{26,27}$ (solid line), overlapping $O_{26,27}$ (dashed line) and reciprocal energy gap $|\Delta_{26,27}|^{-1}$ (dotted line) as a function of the phase $\theta$.
To determine the precise position of the NAPs for the breaking of adiabaticity, we consider the evolution of quantum state as Berry (1984)
@@
where $E_{m}$ is the eigenenergy for instantaneous eigenvector $|\psi_{m}(t)\rangle$. Substituting Eq. 4 into Schrödinger equation yields
@@
With this we define the non-adiabacity $\mathcal{N}_{n}$ for level $n$: $|\psi_{n}(\theta)\rangle$, irrespective of the sweeping rate ${\rm d}\theta/{\rm d}t$, as
$\mathcal{N}_{n}(\theta)=\sum_{m\neq n}N_{n,m}$ (7)
where $N_{n,m}={O_{n,m}}/{\Delta_{n,m}},O_{n,m}={|\langle\psi_{n}(\theta)|\partial_{\theta}|\psi_{m}(\theta)\rangle|}$ and $\Delta_{n,m}={|E_{n}-E_{m}|}$. The larger this value is, the more likely the system will break the adiabaticity, yielding excitation to the bulk state in the pumping process. One can check immediately that for the model in Fig. 1 (c) and Eq. 2, the peaks in $\mathcal{N}_{n}$ happen at $t=T/4$ and $43T/$ 4.
This quantity of non-adiabacity $\mathcal{N}_{n}$ explains our central idea in a much clearer way. In regime A, $\mathcal{N}_{n}(\theta)\sim 1/\sqrt{L}$; and in regime B, $\mathcal{N}_{n}(\theta)\sim\text{constant}$. This is because in regime B, $1/\Delta\sim L$ because both wave functions are delocalized with amplitude $/1\sqrt{L}$, and $O\sim 1/L$, and hence the adiabatic condition is fulfilled by controlling $d\theta/dt$, which is essential for the equivalent dynamics between Eq. 1 and the LZS model. However at the NAPs, this value of non-adiabacity may become very large. This picture is confirmed in Fig. 3(d), in which the peaks of $\mathcal{N}_{n}(\theta)$ are used to determine the position of NAPs.
Figure 4: Distribution of wave function in spacetime and its mean trajectories (green s) for three pump times: (a) $T=1050$, (b) $T=1400$ and (c) $T=9800$ with disorder strength $W=0.08$ and $L=42$. (d) Occupation probability $\rho$ versus pump time $T$ for edge state $|\psi_{26}\rangle$ and bulk state $|\psi_{27}\rangle$. Note that the three pump times are chosen from three extrema of $\rho_{26}$.
@@
ESP with disorder: The LZS interference and the non-adiabacity at the NAPs provide a unique way for us to understand the ESP with disorder, which is inevitably present in experiments Grabarits et al. (2024). In this case the wave function may become localized with localization length $\xi$ Thouless and Kirkpatrick (1981); Lin and Gong (2024)
@@
where $V$ is defined in Eq. 1, $E$ is the eigenvalue and $W$ is the disorder strength. When $\xi\ll L$, where $L$ is the chain length, the state with eigenvalues $E$ should be regarded as localized. In this case, the edge state, if existed, is impossible to be transferred from the left end to the right end, since the overlap between the wave functions, ${O}_{n,m}$, will become exponentially small. In the following, we focus on the case of $\xi\gg L$, where the bulk states are not fully localized. This condition will lead to intriguing physics not observed in the previous experiments. We first present the dynamics for a short chain $L=42$ in Fig. 4, in which the behaviors of these remarkable results are similar to that of a clean chain. Surprisingly, with the increased size $L=105$, the dynamics are totally different. We find that even for a long pumping time $T=10000$, the left edge state cannot be transferred to the right end. The reason is presented in Fig. 5 (b), showing of anticrossing between the edge state and the bulk state at some $\theta$. We find that the edge state will submerge into the bulk states, thus for $T=10^{4}$ and $T=5\times 10^{3}$ [see Fig. 5 (c-d)], the ESP fails to happen. We calculate the non-adiabaticity for the edge state $|\psi_{65}\rangle$ with respect to its nearby excited states $|\psi_{66}\rangle$ and $|\psi_{67}\rangle$ in (e, f) respectively, showing of dramatic enhanced non-adiabaticity at the second NAP due to the anticrossing effect. The presence of anticrossing between the edge state and the bulk states implies a new mechanism for the non-adiabaticity in ESP.

### v2

@@ Abstract
The adiabatic edge state pumping (ESP) in one dimensional model, which has important applications in topological phase transition and quantum simulation, has been widely performed in both theories and experiments. This phenomenon has been observed in some systems with sizes $L=9-100$, and it seems that due to the topological protection, the ESP can be survived even in the presence of weak random potential. Yet the fundamental issues of adiabaticity for this process have not been clarified. In this paper, we revisit this problem and show that this process involves two non-adiabatic points during the transition between the edge state and bulk state, yielding non-abadiatic physics. As a result, the ESP can be described by the Landau-Zener-Stückelberg (LZS) interference process, in which the relative phase between the edge state and the bulk state determine the fate of the edge state during pumping. Furthermore, in a relatively long chain with weak disorder, the ESP can break down due to the anti-crossing of the edge state and the bulk edge states. We unveil these physics in terms of non-adiabaticity. The new mechanisms for ESP unveiled in this work is readily accessible in experiment, and shall therefore offer a down-to-earth platform for the intriguing LZS dynamics in terms of edge states.
@@
Introduction: The discovery of topological matters have stimulated widespread investigations of their edge states ensured by bulk-edge correspondence in quantum simulation Qi and Zhang (2011); Hasan and Kane (2010); Schnyder et al. (2008); Ozawa et al. (2019); Georgescu et al. (2014). In one dimensional (1D) topological models, the edge states are localized at the open ends, which can gradually transfer from one end to the other by adiabatically tuning the parameters, termed as Thouless (or adiabatic) pumping Thouless (1983); Citro and Aidelsburger (2023); Aleiner and Andreev (1998); Nakajima et al. (2016); Sun et al. (2022); Jürgensen et al. (2021, 2023); Cerjan et al. (2020). It has been believed that when the pumping process is slow enough, this kind of adiabatic process can always be achieved. Following this tenet, the adiabatic pumping in a chain with quasiperiodic potentials were experimentally investigated using coupled single-mode waveguides Lahini et al. (2009); Kraus et al. (2012); Rechtsman et al. (2013); Tambasco et al. (2018); Wang et al. (2019); Jürgensen et al. (2021). This idea was also generalized to higher dimensions in terms of synthetic dimensions Yuan et al. (2018); Zilberberg et al. (2018); Yang et al. (2020). Since the edge states are protected by topology, this adiabatic pumping withstands weak disorders Nakajima et al. (2021); Kolodrubetz et al. (2018) and nonlinear interactions Jürgensen et al. (2021); Tangpanitanon et al. (2016); Mostaan et al. (2022); Jürgensen et al. (2023); Qian et al. (2011). In recent years, this adiabatic pumping has been widely examined using electromechanical patches Xia et al. (2021); Wang et al. (2023a); Rosa et al. (2019); Frank et al. (2022), acoustic waveguides Apigo et al. (2019); Ni et al. (2019); Chen et al. (2021); You et al. (2022); Hu et al. (2024), photonic crystals Ke et al. (2016); Meier et al. (2016); Alpeggiani and Kuipers (2019); Cheng et al. (2022); Adiyatullin et al. (2023), and even in the Floquet-Bloch band Upreti et al. (2020); Minguzzi et al. (2022); Adiyatullin et al. (2023). The chain sizes for all systems above range about 9 - 100. The related adiabatic transport of matter in terms of Thouless pumping even with many-body interaction was also explored in experiments Schweizer et al. (2016); Viebahn et al. (2024); You et al. (2022) for applications in quantum matter transport, state transfer and information processing.
@@
This edge state pumping (ESP) is fundamentally different from the adiabatic transition in the two-level system Cohen et al. (2019); Berry (1984), in which the adiabaticity is guaranteed by the finite gap between the two levels. In the continuum limit for many-level system, the number of energy levels is mainly determined by the lattice size $L$, in which the larger the size $L$ is, the smaller the level spacing $\Delta$ will be. Thus there are two competitive ingredients to influence the adiabaticity during pumping: the overlap between the edge state and its nearest neighboring bulk state, and the energy gap between them, both of which decrease with the increasing of $L$. Therefore when the former term decreases faster than the latter, the adiabatic condition can still be fulfilled; otherwise, it becomes possible that the adiabatic condition may break down, yielding non-adiabatic evolution. Therefore, while the ESP has been verified in experiment Kraus et al. (2012); Rechtsman et al. (2013); Switkes et al. (1999); Lahini et al. (2009); Rosa et al. (2019); Frank et al. (2022); Verbin et al. (2015, 2013); Zilberberg et al. (2018); Yang et al. (2020); Nakajima et al. (2021); Kolodrubetz et al. (2018); Jürgensen et al. (2021); Tangpanitanon et al. (2016); Mostaan et al. (2022); You et al. (2022); Jürgensen et al. (2023); Xia et al. (2021); Wang et al. (2023a); Apigo et al. (2019); Ni et al. (2019); Chen et al. (2021); Hu et al. (2024); Ke et al. (2016); Meier et al. (2016); Alpeggiani and Kuipers (2019); Cheng et al. (2022); Adiyatullin et al. (2023); Viebahn et al. (2024); Xiang et al. (2023); Rajagopal et al. (2019); Cerjan et al. (2020), the stringent validity of adiabatic condition, which is essential for ESP, remains yet to be verified. Furthermore, the disorder effect in ESP also needs to be clarified Rajagopal et al. (2019); Yu et al. (2020); Cerjan et al. (2020); Tangpanitanon et al. (2016). These two issues will affect ESP severely in the large $L$ limit with vanished coupling between the edge states Chen et al. (2021).
@@
We then revisit this problem and present a new picture for the ESP in these models. (I) We show that there are two non-adiabatic points (NAPs) during the ESP process, at which the adiabaticity breaks down even in the slowly-varying limit. At the doublet of NAPs, the ESP process reduces to the Landau-Zener-Stückelberg (LZS) interference Berry (2009); Shevchenko et al. (2010); Wang et al. (2015); Boyers et al. (2020); Chen et al. (2021); Qiao et al. (2023), resulting in transient oscillation of pumping efficiency as evolution time $T$ changes. (II) In the presence of weak disorder when the localization length $\xi$ is larger than the system size $L$, the anticrossing of the bulk states and edge state breaks adiabaticity when the edge state is immersed into the bulk states Lahini et al. (2009). In this specific case, the ESP will cease to happen. These new mechanisms can be readily examined in simulating platforms with ultracold atoms Schomerus (2013); Viebahn et al. (2019); Nakajima et al. (2021), optical and acoustic waveguides Xia et al. (2021); Wang et al. (2023a) and electric circuits Li et al. (2023); Mei et al. (2018), etc..
@@
with incommensurate potential $V_{n}=V[1+\lambda\cos(\pi\alpha n+\theta)]$, where $V$ is the hopping strength between neighboring sites, $\alpha$ is the irrational constant and $\hat{c}_{n}$ ($\hat{c}_{n}^{\dagger}$) is the annihilation (creation) operator at site $n$ [see Fig. 1 (a)]. Hereafter, we use $V=8/15$, $\lambda=0.6$, $\alpha=(\sqrt{5}+1)/2$ (parameters from Ref. Kraus et al. (2012)), with uniform disorder $\xi_{n}\in[-1/2,1/2]$. The adiabatic parameter $\theta$ calibrates the process when the edge state transfers from the left end to the right within a finite pump time $T$, and the energy levels in this model versus the phase $\theta$ are presented in Fig. 1 (b), in which the two thick solid blue lines represent the edge modes, respectively, with opposite energies. We focus on the pumping of edge state with $E>0$ in this paper. To this end, we expand the wave function using $|\psi(t)\rangle=\sum_{n}c_{n}(t)|n\rangle$, where $|n\rangle=\hat{c}_{n}^{\dagger}|0\rangle$, and the coefficient $c_{n}$ is determined by Schrödinger equation $i\partial_{t}|\psi(t)\rangle=\hat{H}(t)|\psi(t)\rangle$, with $\theta=\theta(t)$ to be the time-dependent tuning parameter.
@@
Figure 2: Distribution of wave function density $|\psi(x_{i},t)|^{2}$ in spacetime and its mean trajectories $\bar{X}(t)=\sum_{i}x_{i}|\psi(x_{i},t)|^{2}$ (green lines) for three pump times: (a) $T=1300$, (b) $T=1750$ and (c) $T=9500$, with $L=42$ and without disorder. (d) Occupation probability $\rho_{n}$ versus pump time $T$ for edge state $|\psi_{26}\rangle$ and bulk state $|\psi_{27}\rangle$. Note that the three pump times in panels (a-c) are chosen from three extrema on $\rho_{26}$ marked in (d). Inset shows the optimum pump time $T^{*}$ for the first maximum [blue asterisk in (d)] versus the chain length $L$.
@@
Our central idea is that during the ESP, there are two points at which the adiabatic condition is most vulnerable to break. This can be understood as follows. When the edge state is fully localized at the end, the overlap induced by variation of Hamiltonian $\delta H$, which may be defined as $O\sim{|\langle\psi_{\rm e}(\theta)|\delta H|\psi_{\rm b}(\theta)\rangle|}$ with $|\psi_{\rm e}\rangle$ for the edge state and $|\psi_{\rm b}\rangle$ the bulk state respectively, will decay as $1/L$. Meanwhile, the energy difference $\Delta$ between the bulk and edge states will approach a constant; see regime A in Fig. 1 (b). Furthermore, for $\theta$ in regime B in Fig. 1 (b) when the edge states become extended, the overlap between them goes $O\sim 1/L$, while the energy gap between these two states also scales as $\Delta\sim 1/L$. Under these two conditions the the adiabatic condition can be easily fulfilled by controlling the sweeping rate even in the large $L$ limit Grabarits et al. (2024). The adiabaticity may break down between the regimes A and B with the increasing of system size $L$, which are marked by arrows as NAPs.
@@
Thus when the wave functions is restricted to occur between the edge state and the nearest neighboring bulk state and by projecting the wave function to these two states, the dynamics will be reduced to the LZS interference. To gain some insight into this process, let us first consider the following two-level model [Fig. 1 (c)]
@@
where $\hat{\sigma}_{x}$ and $\hat{\sigma}_{z}$ are Pauli matrices, $g$, $A$ are tunneling amplitude and bias amplitude respectively, and $T$ is the total evolution time. The Landau-Zener tunneling happens at $t=T/4$ and $T=3T/4$. In Fig. 1 (d) the dynamics of wave functions is presented, showing that by slightly changing $T$, the final state can be significantly different. The interference depends strongly on the relative phase $\phi$ between the two levels [Fig. 1 (c)], which has been widely verified in experiments Huang et al. (2011); Quintana et al. (2013); Korkusinski et al. (2017); Wang et al. (2023b).
ESP without disorder: We first consider the case without disorder ($W=0$). Initially, we assume the wave function as the edge state $|\psi_{n}(\theta)\rangle$ and define occupation probability
@@
where $|\psi_{n}(\theta)\rangle$ is the $n$-th instantaneous eigenstate with phase $\theta(t)$ and $|\psi(t)\rangle$ is the time-dependent wave function. The calculation results for different $T$ are presented in Fig. 2 (a) - (c) for size $L=42$, and $n=26$ as the index of the edge state. For $T=1300$ and $9500$, the left-end state successfully transfers to the right end. However, for $T=1750$, the state will not evolve to the right end and do so for $T=9500$ in panel (d). These results are in stark contrast to adiabatic theory in which the larger $T$ is, the better the adiabaticity will be Berry (1984). In Fig. 2 (d), the oscillation of $\rho_{26}$ and $\rho_{27}$ versus pump time $T$, is a signature of strong nonadiabaticity before entering the adiabatic limit. Furthermore, our result means $\rho_{26}+\rho_{27}\simeq 1$, indicating that only these two states are important during evolution, which justifies the reduced LZS dynamics of two levels. This oscillation was found numerically in Ref. Longhi (2019), but with its fundamental origin not revealed. The back and forth of the mean trajectories $\bar{X}(t)$ can be readily measured in the currently available experiments. In the inset of (d), we plot the the pump time $T^{*}$ for the first maximum ESP (in blue asterisk) versus system length $L$, showing that $T^{*}\propto L$, which indicates that the longer the chain is, the longer time $T^{*}$ it takes to achieve efficient ESP.
To clarify the LZS dynamics, we plot the evolution of $\rho_{n}$ ($n=26-29$) for $L=42$ for three typical $T$ in Fig. 3 (a) - (c). A common feature is that the tunnelling occurs only at the two NAPs. In panel (a) at the first NAP, $\rho_{26}$ drops suddenly to about 0.5, lending the other half portion to the bulk $\rho_{27}$. However at the second NAP, the edge state $|\psi_{26}\rangle$ will then reclaim the occupation from the bulk $|\psi_{27}\rangle$, reaching the unity probability. This is a surprising example to realise a perfect ESP yet with non-adiabatic process. In (b) with an even-larger pump time $T=1750$, the edge state finally transfers to the bulk after two tunnellings at the NAPs, with $\rho_{26}\sim 0$. In (c) when $T$ is large enough, the ESP is achieved in the adiabatic limit. This dependence of occupation probabilities $\rho_{n}$ with pump time $T$ corroborates the LZS process shown in Fig. 1 (d).
Figure 3: Evolution of $\rho(t)$ for the edge state $|\psi_{26}\rangle$ and the three nearest bulk states $|\psi_{27}\rangle$, $|\psi_{28}\rangle$ and $|\psi_{29}\rangle$ for pump time (a) $T=1300$, (b) $T=1750$ and (c) $T=9500$ (Here, $L=42$). (d) Non-adiabaticity ${N}_{26,27}$ (solid line), overlapping $O_{26,27}$ (dashed line) and reciprocal gap $|\Delta_{26,27}|^{-1}$ (dotted line) as a function of the phase $\theta$.
In order to determine the precise positions of the NAPs for the breaking of adiabaticity, we consider the evolution of quantum state as Berry (1984)
@@
where $E_{m}$ is the eigenenergy for instantaneous eigenvector $|\psi_{m}(t)\rangle$. Substituting Eq. (4) into Schrödinger equation yields
@@
With this we define the non-adiabaticity $\mathcal{N}_{n}$ for $|\psi_{n}(\theta)\rangle$, irrespective of the sweeping rate ${\rm d}\theta/{\rm d}t$, as
$\mathcal{N}_{n}(\theta)=\sum_{m\neq n}N_{n,m}(\theta)$ (7)
where $N_{n,m}={O_{n,m}}/{\Delta_{n,m}},O_{n,m}={|\langle\psi_{n}(\theta)|\partial_{\theta}|\psi_{m}(\theta)\rangle|}$ and $\Delta_{n,m}={|E_{n}-E_{m}|}$. The larger this value $\mathcal{N}_{n}$ is, the more likely the edge state $|\psi_{n}(\theta)\rangle$ will break the adiabaticity, yielding excitation to the bulk state $|\psi_{m}(\theta)\rangle$ in the pumping process. One can check immediately that for the LZS model in Fig. 1 (c) and Eq. (2), the non-adiabaticity $\mathcal{N}_{n}$ peaks at $t=T/4$ and $43T/$ 4.
Thus the non-adiabaticity $\mathcal{N}_{n}$ encapsulates our central idea in a much clearer way. In regime A, $\mathcal{N}_{n}(\theta)\sim 1/\sqrt{L}$; and in regime B, $\mathcal{N}_{n}(\theta)\sim\text{constant}$. This is because in regime B, the level spacing goes as $1/\Delta\sim L$ and the overlap $O_{n,m}\sim 1/L$ since both wave functions are delocalized with amplitude $\sim 1/\sqrt{L}$. Thus the adiabatic condition is fulfilled by controlling the sweeping rate $d\theta/dt$, which is essential for the equivalent dynamics between Eq. (1) and the LZS model. Moreover specifically at the NAPs, the values of non-adiabacity become very large. This picture is confirmed in Fig. 3(d), in which the peaks of $\mathcal{N}_{26}(\theta)\simeq N_{26,27}$ are used to pinpoint the NAPs.
Figure 4: Distribution of wave function in spacetime and its mean trajectories (green curves) for three pump times: (a) $T=1050$, (b) $T=1400$ and (c) $T=9800$ with disorder strength $W=0.08$ and $L=42$. (d) Occupation probability $\rho$ versus pump time $T$ for edge state $|\psi_{26}\rangle$ and bulk state $|\psi_{27}\rangle$. Note that the three pump times are chosen from three extrema of $\rho_{26}$, whose occupation probabilities are marked in (d).
@@
ESP with disorder: The LZS interference and the non-adiabaticity at NAPs provide a unique way for us to understand the ESP with disorder, which is inevitably present in experiments Grabarits et al. (2024). In this case the wave function may become localized with localization length $\xi$ given by Thouless and Kirkpatrick (1981); Lin and Gong (2024)
@@
where $V$ is the hopping strength, $E$ is the eigenvalue and $W$ is the disorder strength, all defined in Eq. (1). When $\xi\ll L$ with $L$ the chain length, the state with eigenvalues $E$ should be regarded as localized. In this case, the edge state, if existed, is impossible to be transferred from the left end to the right end, since the overlap between the wave functions, ${O}_{n,m}$, will become exponentially small. In the following, we focus on the case of $\xi\gg L$, where the bulk states are not fully localized. This condition will lead to intriguing physics not observed in the previous experiments. We first present the dynamics for a short chain $L=42$ in Fig. 4, in which the behaviors of the results are similar to that of a clean chain in Fig. 2. Surprisingly, with the size increased to $L=105$, the dynamics is totally different. We find that even for a long pump time $T=10000$, the left edge state cannot be transferred to the right end. The reason is presented in Fig. 5 (b), showing the anticrossing between the edge state and the bulk state at some $\theta$. We find that the edge state will gradually submerge into the bulk, and thus the ESP fails to happen for $T=10^{4}$ and $T=5\times 10^{3}$ [see Fig. 5 (c-d)]. This failure is attributed to the non-adiabaticities for the edge state $|\psi_{65}\rangle$ with respect to its adjacent excited states $|\psi_{66}\rangle$ and $|\psi_{67}\rangle$: $N_{65,66}$ and $N_{65,67}$ in panels (e, f) respectively, which peak the most at the second NAP due to their anticrossing. The presence of anticrossing between the edge and bulk states implies a new mechanism for the non-adiabaticity in ESP.